CARS AND INFRARED STUDIES OF ^{32}S^{16}O_{3}: HOW CAN SUCH A SIMPLE MOLECULE GIVE SUCH A COMPLEX SPECTRUM?

\underline{JEFFREY BARBER}, ENGELENE t.H. CHRYSOSTOM, TONY MASIELLO, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, Corvallis, OR 97331; ARTHUR MAKI, 15012 24^{th} Ave. S. E. Mill Creek, WA 98012; ALFONS WEBER, National Science Foundation, Arlington, VA 22230 and National Institute of Standards and Technology, Gaithersburg, MD 20899; THOMAS A. BLAKE, ROBERT L. SAMS, Pacific Northwest National Laboratory, Richland, WA 99352.

Sulfur trioxide is an important participant in reactions in the upper atmosphere and also in a number of industrial processes. It is a D_{3h} planar oblate top whose spectroscopy is surprisingly incomplete, in part due to its extreme corrosive properties. We report here a comprehensive investigation of all four fundamental bands of ^{32}S^{16}O_{3}. High resolution (0.001 cm^{−1}) coherent anti-Stokes Raman scattering (CARS) has been used at Oregon State University to examine the Q-branch structure of the IR-inactive ν_{1} symmetric stretching mode of ^{32}S^{16}O_{3}. This spectrum reveals two intense Qbranch regions, with complex vibrational-rotational structure. The modeling of this has been challenging and involves a subtle combination of Fermi-resonances, indirect Coriolis interactions, and l-resonances with nearby hidden states; 2ν_{4} (l= 0, ± 2), ν_{2} + ν_{4} (l = ± 1), 2ν_{2} (l =0). The analysis of the perturbed ν_{1} spectrum, achieved by locating these states via high-resolution infrared hot-band studies performed at Pacific Northwest National Laboratory, will be presented in this talk.